OpenStudy (anonymous):

-4 sqrt x+9=20 need help please!

OpenStudy (anonymous):

is it \(-4\sqrt{x+9}=20\) or is it \[-4\sqrt{x}+9=20\]?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

then there is no answer, because the left hand side \(-4\sqrt{x+9}\) is negative, but the right hand side \(20\) is positive

OpenStudy (anonymous):

\(\sqrt{a}\) is always greater than or equal to zero, so when you multiply by \(-4\) the term will be less than or equal to zero, and it cannot equal a positive number

OpenStudy (anonymous):

so there is no answer?

OpenStudy (anonymous):

No, there is no solution like satellite73 says....

OpenStudy (anonymous):

yes, there is no solution to this

OpenStudy (anonymous):

thank you!